OpenStudy (anonymous):

-5+1/8a=6 what is the value of a

OpenStudy (anonymous):

(-5 + 1) whole in the numerator or 1 is the numerator??

OpenStudy (anonymous):

1 ove 8

OpenStudy (anonymous):

ops i ment over

OpenStudy (anonymous):

-5 is separate??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is it? \[-5 + \frac{1}{8a} = 6\] Right??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait it a 5 not an 8

OpenStudy (anonymous):

Multiply by 5a both the sides: \[-25a + 1 = 30a \implies 1 = 55a \implies \color{green}{ a = \frac{1}{55}}\]

OpenStudy (anonymous):

Is this the question? \[-5 + \frac{1}{5a} = 6\]

OpenStudy (anonymous):

yes srry bout the confusion O_O

OpenStudy (anonymous):

It is okay.. a = 1/55 is the answer..

OpenStudy (anonymous):

thx